# Evaluating Human-AI Collaboration: A Review and Methodological Framework

### v1

@@ 3.1 Advancements in Evaluation Frameworks
3.2 Advancements in Evaluation Tools and Methodologies: Enabling Real-time Insights and Adaptive Collaboration
Technological advancements are transforming the landscape of HAIC evaluation, enabling real-time monitoring, adaptive collaboration, and deeper insights into human-AI interactions.
AI-powered analytics and real-time monitoring tools, discussed by Wang et al. (2019) [77], are enhancing the precision and responsiveness of HAIC evaluations. These tools can automatically track key performance indicators (KPIs) such as completion time, error rates, and user interactions, providing immediate feedback on collaboration efficacy and enabling timely interventions for improvement.
Advancements in interactive machine learning (iML) interfaces, explored by Saha et al. (2023) [78], are fostering grater transparency and user understanding. By providing explanations for AI-generated recommendations and allowing users to adjust system parameters, iML interfaces can improve user comprehension, trust, and satisfaction, ultimately contributing to a more effective and collaborative partnership between humans and AI. Saha et al. [78] evaluate the effectiveness of iML interfaces using a combination of user surveys, interviews, and task-based performance metrics.
3.3 Analysis of Existing Approaches
@@
Sharma et al. (2023) [49] look at how conversational empathy improves in mental health applications using qualitative measures. Reddy et al. (2021) [48] suggest a qualitative framework for evaluating the use of AI in healthcare, with a focus on ethical issues and user perspectives. Their framework includes structured interviews with healthcare providers and patients, and thematic analysis to identify key concerns such as data privacy, trust in AI diagnoses, and the perceived impact on the patient-provider relationship. Their study underscores the necessity for AI systems to align with ethical standards and user expectations to be effectively integrated into healthcare practices.
By capturing the nuanced perspectives of users and stakeholders, these qualitative studies offer a rich understanding of the human factors that shape the adoption and impact of HAIC. However, the subjective nature of qualitative data can pose challenges in terms of generalizability and require substantial resources for data collection and analysis.
The Mixed-Methods Promise: Towards a Holistic and Contextualized Evaluation
The limitations of purely quantitative or qualitative approaches have spurred the adoption of mixed-methods evaluations, aiming to capture both objective performance and subjective experiences in a comprehensive manner.
@@
In the realm of design and art, Arias-Rosales (2022) [63] combines quantitative shape generation metrics with qualitative assessments of perceived value by designers. Their study uses quantitative metrics to measure the efficiency and novelty of AI-generated shapes, while also conducting interviews and surveys with designers to gather insights on the aesthetic value and usability of these shapes. This approach revealed that while AI-generated designs were technically innovative, the perceived value varied significantly among designers, highlighting the importance of subjective aesthetic judgment in evaluating creative outputs.
Similarly, in finance, Chakravorti et al. (2022) [66] integrate quantitative detection rates with qualitative assessments of trust and perception to evaluate the efficacy of AI prediction markets. They employ quantitative methods to analyze the accuracy and reliability of AI predictions in financial markets, alongside qualitative interviews and surveys with traders and financial analysts. The study found that despite high detection rates, the trust in AI predictions varied, with qualitative data uncovering concerns about the transparency and interpretability of AI algorithms.
@@
Yang et al. (2022) [55] look into how AI can model human behavior using both quantitative and qualitative data. They use quantitative metrics such as task completion rates and error frequencies to evaluate AI performance, while also gathering qualitative data through user interviews and behavioral observations. Vössing et al. (2022) [79] study includes quantitative measures of system accuracy and user performance metrics, coupled with qualitative trust assessments obtained from user interviews and surveys. The combination of these methods revealed that transparency significantly impacts user trust and acceptance, even when the AI system’s performance is objectively high.
Huang & Rust (2022) [67] propose a mixed-methods framework for collaborative AI in marketing, acknowledging the importance of both performance and ethical considerations.These mixed-methods techniques provide the possibility for a comprehensive and contextual evaluation of HAIC, recognizing the intricate relationship between objective measurements and subjective perceptions. They employ quantitative analytics to measure marketing campaign performance metrics, such as click-through rates and conversion rates, while also using qualitative methods like focus groups and in-depth interviews to assess ethical concerns and user satisfaction. This dual approach provides a comprehensive evaluation of the AI’s effectiveness and its alignment with ethical standards.
3.4 The domain-specific challenges necessitate a tailored evaluation
The application domains for HAIC are diverse, each presenting unique challenges that necessitate tailored evaluation approaches. These challenges stem not only from the technical complexities of integrating AI into different workflows, but also from the varying priorities and expectations of stakeholders in different sectors. As a result, the evaluation methods employed across HAIC research vary significantly, reflecting the need to adapt assessment strategies to the specific context of each domain.
@@ 3.4 The domain-specific challenges necessitate a tailored evaluation
In the healthcare sector, ensuring diagnostic accuracy and patient safety is paramount. To find out how well AI-assisted tools work in oncology and radiology, studies like Vollmuth et al. (2023) [41] and Van Leeuwen et al. (2022) [42] use numbers like sensitivity, specificity, and outcome improvements. However, the high-stakes nature of medical decisions and the potential for unintended consequences necessitate additional layers of evaluation. Qualitative approaches like interviews and focus groups (Farić et al., 2024 [14]; Calisto et al., 2022 [45]) can capture user perceptions, identify potential biases (Timmons et al., 2023 [43]), and ensure that AI systems are implemented in a way that aligns with ethical principles and patient values (Reddy et al., 2021 [48]).
@@
In the creative arts, HAIC evaluation often requires a different approach. While quantitative metrics can measure the technical performance of AI systems, such as the speed or accuracy of generating artistic outputs, qualitative methods such as interviews, user feedback, and case studies are crucial for comprehending the influence of AI on creativity, collaboration, and the overall design process (Rezwana et al., 2023 [46]; Verheijden & Funk, 2023 [57]). Additionally, ethical considerations, such as the potential for AI to perpetuate biases or undermine the role of human creativity, necessitate careful evaluation through qualitative assessments and critical analysis. Moreover, the cooperative aspect of HAIC plays a crucial role in enhancing human creativity, since AI’s data-driven insights stimulate inventive solutions [80].
@@
The financial sector employs a mix of quantitative and qualitative methods to address HAIC’s unique challenges in this domain. Quantitative metrics like detection rates and false positives are crucial for evaluating the effectiveness of AI in fraud detection (Al-Fatlawi et al., 2023 [54]). However, understanding the impact of AI on decision-making processes, investor trust, and market stability requires qualitative insights through surveys, interviews, and case studies (Chakravorti et al., 2022 [66]). Mixed-methods approaches, which combine quantitative and qualitative data, offer a more comprehensive understanding of the complex interactions between humans, AI, and financial markets.
In summary, the review of evaluation approaches demonstrates significant variation in collaboration methods, domain-specific challenges, and methodological approaches. However, one significant obstacle remains: the lack of a comprehensive, mixed-methods evaluation framework that covers a wide range of industries. This gap restricts our ability to completely evaluate HAIC’s effectiveness and maximize its potential across domains.
The Path Forward: A Unified Evaluation Framework to Bridge the Gap in HAIC Assessment
The analysis of existing approaches to HAIC evaluation reveals a landscape marked by diversity and fragmentation. Quantitative, qualitative, and mixed-methods approaches all give useful information, but the lack of a single, flexible framework makes it harder to get a full picture of how well HAIC works and what it can do in different fields.
This gap is particularly evident when considering the multifaceted nature of HAIC. Quantitative metrics, while essential for measuring objective performance, often fail to capture the nuances of human-AI interaction dynamics, such as trust, satisfaction, and ethical considerations. Qualitative methods, on the other hand, can provide rich insights into these subjective experiences but may lack the rigor and generalizability of quantitative data.
Therefore, it is imperative to develop a comprehensive, mixed-methods evaluation framework. This framework should be adaptable across a broad spectrum of industries, accommodating the unique characteristics and challenges of each while providing a cohesive approach to assessing the effectiveness, challenges, and opportunities of HAIC. By integrating objective performance metrics with subjective experiences, such a framework can offer a more nuanced understanding of how HAIC systems perform in real-world contexts and how they impact various stakeholders.
This unified framework would not only address the current methodological fragmentation but also facilitate cross-study comparison, enhance the generalizability of findings, and promote standardized evaluation practices. It would enable researchers, practitioners, and policymakers to make informed decisions based on a more holistic understanding of the complex interactions between humans and AI.
In healthcare, this framework could integrate measures of diagnostic accuracy with assessments of user trust and ethical considerations, ensuring that AI-assisted tools are both effective and aligned with patient values. In creative fields, it could incorporate quantitative metrics of creative output alongside qualitative assessments of the impact on artistic processes and collaboration. In finance, it could combine quantitative measures of risk and return with qualitative assessments of trust, transparency, and the impact on decision-making.
By addressing the limitations of existing approaches and providing a standardized yet adaptable framework, we can unlock the full potential of HAIC across diverse domains. This would enable us to harness the power of AI in ways that augment human capabilities, foster innovation, and ultimately contribute to societal well-being. The development of such a framework is not merely an academic pursuit but a critical step towards ensuring the responsible and impactful integration of AI into human workflows.
@@ 4 A Structured Framework for Human-AI Collaboration
4.1 Factors of Identifying the Human-AI Collaboration
Based on the previous discussion (Section 2) the evaluation of HAIC is centered around three primary factors: Goals, Interaction, and Task Allocation. These factors are critical for understanding and enhancing HAIC systems. While tasks themselves are fundamental to HAIC, they are considered the context within which collaboration occurs, rather than a core factor of the evaluation framework. The nature of the task will influence how the other three factors are approached, but the tasks themselves are not directly evaluated by this framework.The literature highlights the importance of these factors in evaluating HAIC.
Research by Li et al. [81] and Vossing et al. [65] underscores the significance of evaluating goals through metrics such as the learning curve and prediction accuracy. Studies by Hinsen et al. [82] and Amershi et al. [83] emphasize the need for user friendly communication methods and clear interaction mechanisms for successful HAIC. Additionally, Xiong et al. [84] stress the importance of measuring query efficiency and expertise utilization for effective task allocation.
Following the identification of these primary factors, we further explore the subfactors and metrics that can be used to evaluate each factor. These subfactors and metrics are designed to provide a comprehensive understanding of the effectiveness, efficiency, and usability of HAIC systems. The structured framework presented here serves as a guide for researchers, practitioners, and policymakers to evaluate and optimize HAIC systems across various domains and applications.
@@ 4.2 Subfactors & Metrics of the Framework
To promote further analysis and evaluation of the different types of interactions in HAIC, the three primary factors are further analyzed into measurable subfactors, which can be evaluated using a set of metrics. These are presented in Table 2. Following we provide a detailed description of the subfactors and metrics associated with each primary factor. Please note that Table 3 summarizes all metrics discussed in this section, including their formal definition.
{adjustwidth}
-1cm-1cm Primary Factor Subfactor Human-AI Collaboration Metrics for Evaluation Goals Individual Goals AI’s Goal: Learn Efficiently, Increase Accuracy. Human’s Goal: Provide Effective Teaching, Achieve Task Objectives. * Learning Curve (rate of improvement over time), * Prediction Accuracy (percentage of correct predictions), * Teaching Efficiency (time/resources spent per unit of learning) [81, 85, 86]. Collective Goals Shared Goal: Maximizing the overall performance of the AI system. * Overall System Accuracy (percentage of correct outputs), * Objective Fulfillment Rate (percentage of tasks successfully completed) [65, 87]. Interaction Communication Methods Query and Response Mechanism. * Ease of Use (user satisfaction with communication methods), * Clarity of Communication (understandability of queries and responses) [82, 83] Feedback Mechanisms Human feedback on AI queries, AI adjustments based on feedback. * Frequency of Feedback (number of feedback instances), * Feedback Quality (relevance and specificity of feedback), * Feedback Impact (measured change in AI performance due to feedback) [83, 88, 82] Adaptability AI’s ability to adapt query strategy and learning based on human input. * Strategy Adaptation Effectiveness (improvement in query selection based on feedback), * Adaptability Score (composite metric of flexibility and responsiveness) [82, 89, 90] Trust and Safety Trust in AI system, ensuring user safety. * Trust Score (level of user trust in AI system) [91], * Safety Incidents (number of safety-related issues) [92], * Error Reduction Rate (decrease in errors due to AI assistance) [4], * Confidence (level of user trust in AI recommendations) [93] Task Allocation Complementarity AI handles data analysis and uncertainty estimation; Humans provide expertise and clarification. * Query Efficiency (amount of information gained per query), * Expertise Utilization (degree to which human expertise is leveraged) [84] Flexibility Adjusting query strategy based on interaction dynamics. * Adaptability Score (same as in Interaction), * Dynamic Task Allocation (ability to adjust task distribution based on real-time needs) [3] Efficiency Streamlining the learning process through targeted queries. * Model Improvement Rate (rate of improvement in AI performance) [65, 84], * Learning Curve (same as in Goals), * Resource Utilization (efficiency in using computational resources) [65] Responsiveness AI’s ability to rapidly adjust to new inputs and feedback. * Response Time (time taken to respond to queries or feedback), * Real-time Performance (ability to handle time-sensitive tasks) [82, 94] Collaborative Decision Making Joint decisions on data labeling and correction of AI’s model. * Impact of Corrections (measured improvement in system performance through human intervention) [95, 96], * Decision Effectiveness (impact of decisions on overall task success) [97, 98] Continuous Learning AI’s ability to learn from human input and adjust its model. * Model Improvement Rate (same as in Efficiency), * Learning Curve (same as in Goals), * Knowledge Retention (ability to retain learned information over time) [99] Mutual Support AI’s ability to provide support to humans in the learning process; Humans support by providing information. * Error Reduction Rate (decrease in errors due to AI assistance) [4], * Confidence (level of user trust in AI recommendations) [100], * Task Completion Time (with and without AI support) [101] Robustness Ensuring AI system’s robustness and reliability. * Adversarial Robustness (resilience to adversarial attacks) [102], * System Reliability (overall system uptime and fault tolerance) [103]
@@
Goals
@@
The Goals factor ensures that the collaboration has a clear direction and that both human and AI efforts are aligned towards shared outcomes. This factor encompasses two key subfactors:
  • Individual Goals: The specific aims of each participant. For the AI, this could be learning efficiently or maximizing accuracy. For humans, it might involve providing effective teaching or achieving task objectives. Misalignment of individual goals can hinder collaboration, as highlighted in studies on migrating to HAIC, which emphasize the importance of goal congruence [40].
@@
    • Learning Curve: Measures the rate of improvement over time, indicating the effectiveness of human teaching and AI learning. This is plotted by graphing the performance of the AI (like accuracy) against the number of training samples or training iterations. A steep learning curve suggests the AI is learning effectively [81].
@@
    • Prediction Accuracy: The proportion of correct predictions made by the AI model out of all predictions. This can be measured using metrics like accuracy, F1 score, or area under the ROC curve, depending on the nature of the task [86].
@@
    • Teaching Efficiency: The gain in AI performance per unit of teaching effort by the human. This could be measured by tracking improvements in the AI’s accuracy or error rate after receiving human input. This concept is explored in works like that of Cakmak and Thomaz [85]
  • Collective Goals: Shared objectives that maximize the overall performance of the HAIC system. For example, in a medical diagnosis system, the collective goal might be to improve patient outcomes through accurate and timely diagnosis. These collective goals are often broader than individual goals and necessitate the combined efforts of both humans and AI.
@@
    • Overall System Accuracy: It can be calculated as a weighted average of different performance metrics, reflecting the relative importance of each task in achieving the overall goal. Evaluates the general accuracy of the AI-human collaborative system, tailored to the specific objectives of the system. [65].
@@
    • Objective Fulfillment Rate: The percentage of predefined goals that are successfully achieved within the HAIC collaboration. This metric is often used in project management contexts to track progress towards overall objectives and can be adapted to assess the success of HAIC collaborations [87].
@@
The Interaction factor underscores the critical communication mechanisms through which humans and AI systems exchange information, provide mutual feedback, and adaptively respond to each other’s inputs. Key subfactors within the Interaction factor include:
  • Communication Methods: This involves evaluating the clarity, accessibility, and intuitiveness of the interfaces and protocols through which humans and AI interact. For instance, the clarity of communication in a medical diagnosis system might be assessed by how well the AI can explain its reasoning to the doctor in an understandable way [105, 90]. Effective communication methods ensure efficient information exchange and mutual understanding, which can be evaluated using metrics such as:
@@
    • Clarity of Communication: Can be assessed through qualitative methods like expert review or user feedback, or through quantitative measures like the proportion of correctly interpreted queries or the time taken to resolve ambiguities.
@@
    • Ease of Use: Often measured through Likert scale questionnaires or user interviews to gauge the user’s subjective experience with the system.
  • Feedback Mechanisms: These mechanisms assess how input from human operators is used to guide AI learning and behavior adjustments. For instance, in a customer service chatbot, feedback mechanisms could involve analyzing how user ratings or corrections influence the chatbot’s future responses. It’s crucial that AI systems can not only receive but also meaningfully integrate feedback to enhance their interactions with humans [83, 88]. The effectiveness of feedback mechanisms can be measured through:
    • Frequency of Feedback (number of feedback instances): This is a simple count of how often users provide feedback, indicating engagement and potentiallearning opportunities for the AI, as highlighted in Amershi et al.’s (2014) [83] work on power to the people.
@@
    • Feedback Quality (relevance and specificity of feedback): This can be measured qualitatively through expert review or quantitatively by analyzing the correlation between feedback and subsequent AI performance improvements as suggested by Dautenhahn (2007) [88] in the context of socially intelligent robots.
@@
    • Feedback Impact (measured change in AI performance due to feedback): This quantifies the tangible effect of feedback on AI behavior or output quality, often measured by improvements in accuracy, precision, or other relevant metrics.
  • Adaptability: This measures the AI’s ability to modify its behaviors based on feedback and changes in the environment or task conditions. Adaptability is essential for AI systems to remain effective as collaboration needs evolve [106, 107]. For example, in an autonomous vehicle, the AI system should be able to adapt its driving style based on the driver’s feedback or changes in traffic conditions. Adaptability can be evaluated through:
    • Adaptability Score (composite metric of flexibility and responsiveness): This metric combines measures of how quickly and effectively the AI responds to changes and how flexibly it adjusts its strategies, which can be derived from studies like that of El-Assady et al. (2022) [90] on communication biases in HAIC.
@@
    • Strategy Adaptation Effectiveness (improvement in query selection based on feedback): In systems that involve querying for information, this metric measures whether the AI learns to ask more effective questions based on human responses, potentially incorporating insights from research on semantic interaction challenges by Wenskovitch et al. (2020) [89].
  • Trust and Safety: Ensuring trust in the AI system and the safety of users is critical for effective HAIC. Metrics to evaluate trust and safety include:
@@
    • Trust Score: The level of user trust in the AI system, often measured through surveys and trust scales [100].
@@
    • Safety Incidents: The number of safety-related issues encountered, used to assess the reliability and safety of the AI system.

### v2

@@ 2.3.1 Advancements in Tools, Methodologies, and Real-Time Collaboration for HAIC
Technological advancements further enable real-time monitoring and adaptive collaboration in HAIC. Recent studies, such as Woelfle et al. (2024) [11], emphasize the importance of benchmarking tools that evaluate human-AI collaboration through diverse performance metrics. These metrics extend beyond traditional KPIs like task completion time and error rates, incorporating dimensions such as evidence appraisal quality and the adaptability of AI systems to human inputs. Such tools provide actionable feedback on collaboration efficacy, fostering dynamic adaptability and enhancing the overall effectiveness of HAIC systems.
Interactive machine learning (iML) interfaces, explored by Saha et al. (2023) [72], represent a significant leap in fostering transparency and user engagement. By providing explanations for AI-generated recommendations and allowing users to adjust system parameters, iML interfaces enhance user comprehension, trust, and satisfaction. Saha et al. evaluate these interfaces using a combination of user surveys, interviews, and task-based performance metrics, demonstrating their potential to facilitate effective human-AI partnerships.
Together, these advancements underscore the importance of integrating ethical considerations, real-time analytics, and interactive tools to develop comprehensive evaluation frameworks for HAIC. These tools and methodologies not only enhance our understanding of human-AI collaboration but also pave the way for more adaptive, transparent, and ethical partnerships between humans and AI systems.
2.3.2 Analysis of Existing Approaches
@@
Sharma et al. (2023) [43] look at how conversational empathy improves in mental health applications using qualitative measures. Reddy et al. (2021) [42] suggest a qualitative framework to evaluate the use of AI in healthcare, with a focus on ethical issues and user perspectives. Their framework includes structured interviews with healthcare providers and patients, and thematic analysis to identify key concerns such as data privacy, trust in AI diagnoses, and the perceived impact on the patient-provider relationship. Their study underscores the need for AI systems to align with ethical standards and user expectations to be effectively integrated into healthcare practices.
By capturing the nuanced perspectives of users and stakeholders, these qualitative studies offer a rich understanding of the human factors that shape the adoption and impact of HAIC. However, the subjective nature of qualitative data can pose challenges in terms of generalizability and requires substantial resources for data collection and analysis.
2.3.3 The Mixed-Methods Promise: Towards a Holistic and Contextualized Evaluation
The limitations of relying solely on quantitative or qualitative approaches have led to the adoption of mixed-methods evaluations, which integrate objective performance metrics with subjective experiences. This approach provides the flexibility needed to tailor evaluations to specific domains while balancing the rigor of quantitative data with the depth of qualitative insights.
@@ 2.3.3 The Mixed-Methods Promise: Towards a Holistic and Contextualized Evaluation
In creative domains such as design and art, Arias-Rosales (2022) [57] combined quantitative shape generation metrics with qualitative assessments of designers’ perceived value. This dual approach highlighted that while AI-generated designs were technically innovative, their subjective aesthetic value varied significantly among designers, underlining the critical role of qualitative judgment in evaluating creative outputs. Similarly, in finance, Chakravorti et al. (2022) [60] evaluated AI prediction markets by pairing detection rates with trader and analyst perceptions. Their mixed-methods approach revealed that high prediction accuracy did not necessarily translate into trust, exposing concerns about transparency and interpretability.
@@
Mixed methods also demonstrate utility across domains. For instance, Yang et al. (2022) [49] modeled human behavior by combining task completion rates and error frequencies with user interviews, while Vössing et al. (2022) [73] integrated system accuracy metrics with trust assessments, uncovering the significant impact of transparency on user acceptance. In marketing, Huang & Rust (2022) [61] employed a mixed-methods framework that combined performance metrics, such as click-through rates, with ethical and user satisfaction insights. Together, these studies illustrate how mixed methods enable comprehensive evaluations that capture both technical performance and contextual user perceptions.
2.3.4 Tailored Evaluation for Domain-Specific Challenges
HAIC’s diverse application domains present unique challenges that demand tailored evaluation approaches, reflecting both technical complexities and the distinct priorities of stakeholders. As discussed in the Appendix A of this work, the evaluation of creative and linguistic AI requires tailored methodologies due to its inherent complexities.
@@ 2.3.4 Tailored Evaluation for Domain-Specific Challenges
In the healthcare, diagnostic accuracy and patient safety are critical. Studies like Vollmuth et al. (2023) [35] and Van Leeuwen et al. (2022) [36] use metrics like sensitivity, specificity but the high-stakes nature of medical decisions require additional qualitative approaches like interviews and focus groups (Farić et al., 2024 [74]; Calisto et al., 2022 [39]) to capture user perceptions, identify biases (Timmons et al., 2023 [37]), and ensure alignment with ethical principles (Reddy et al., 2021 [42]).
@@
In creative arts, quantitative metrics (e.g. speed, output quality) are complemented byqualitative methods such as interviews and case studies (Rezwana et al., 2023 [40]; Verheijden & Funk, 2023 [51]) to asses AI’s influence on creativity and collaboration. Ethical concerns regarding bias and the potential undermining of human creativity, underscore the need for a balanced evaluation [75].
@@
In finance quantitative metrics such as detection rates and false positives are crucial for fraud detection (Al-Fatlawi et al., 2023 [48]). However, understanding AI’s impact on decision-making and market stability also requires qualitative insights through surveys and interviews (Chakravorti et al., 2022 [60]). Mixed-methods approaches, provide a comprehensive view of human-AI interaction in financial markets. For a detailed discussion on the challenges specific to creative and linguistic AI, please refer to Appendix A.
In summary, while evaluation approaches vary significantly across domains, a comprehensive, mixed-methods framework is still needed to sully assess HAIC’s effectiveness and potential.
2.4 Towards a Unified Evaluation Framework
While existing HAIC evaluation approaches offer valuable insights, they remain fragmented. A unified framework is needed to integrate objective performance metrics with subjective experiences across diverse domains. Such a framework would support cross-study comparison, enhance generalizability, and standardize evaluation practices.
This framework should adapt to different industries, ensuring AI-assisted tools are both effective and aligned with human needs. In healthcare, it could integrate diagnostic accuracy with user trust; in creative fields, it could assess AI’s impact on artistic processes; and in finance, it could combine risk metrics with transparency evaluations. By addressing methodological gaps, this approach will help harness AI’s potential while ensuring its responsible integration into human workflows.
@@ 3 A Structured Framework for Human-AI Collaboration
Unlike previous studies that focus on isolated aspects of HAIC evaluation, the proposed framework provides a structured approach by setting measurable goals, clarifying tasks and reporting metrics [bolman2013reframing], that integrates multiple evaluation dimensions, organized in three primary factors and specialized subfactors. The primary factors are:
  • Goals – Evaluates individual and collective objectives within HAIC systems, with key metrics such as learning curves and prediction accuracy [76, 59].
  • Interaction – Assesses communication mechanisms, feedback processes, adaptability, and trust. Effective communication and transparency are critical for ensuring human-AI trust and usability [77].
  • Task Allocation – Examines role distribution, efficiency, and decision-making processes, ensuring that AI contributions complement human expertise in dynamic environments [78].
Each factor is further divided into measurable subfactors and associated metrics, allowing for a robust quantitative and qualitative assessment of HAIC effectiveness. To enhance practical applicability, we develop a decision-tree-based evaluation approach. This structured method assists practitioners in selecting relevant evaluation criteria based on the nature of the HAIC system. The decision tree ensures flexibility, allowing for tailored assessments in various domains, including healthcare, finance, manufacturing, and education.
By offering a holistic assessment tool, this framework addresses the key gaps identified in the literature and establishes a standardized methodology for evaluating HAIC in both research and industry applications.
@@ 3.1 Subfactors & Metrics of the Framework
Table 2 summarizes the detailed breakdown of factors to subfactors and their associated metrics.
@@
Primary Factor Subfactor Human-AI Collaboration Metrics for Evaluation
Goals Individual Goals AI’s Goal: Learn Efficiently, Increase Accuracy. Human’s Goal: Provide Effective Teaching, Achieve Task Objectives. * Learning Curve (rate of improvement over time), * Prediction Accuracy (percentage of correct predictions), * Teaching Efficiency (time/resources spent per unit of learning) [76, 79, 80].
Collective Goals Shared Goal: Maximizing the overall performance of the AI system. * Overall System Accuracy (percentage of correct outputs), * Objective Fulfillment Rate (percentage of tasks successfully completed) [59, 81].
Interaction Communication Methods Query and Response Mechanism. * Ease of Use (user satisfaction with communication methods), * Clarity of Communication (understandability of queries and responses) [82, 83]
Feedback Mechanisms Human feedback on AI queries, AI adjustments based on feedback. * Frequency of Feedback (number of feedback instances), * Feedback Quality (relevance and specificity of feedback), * Feedback Impact (measured change in AI performance due to feedback) [83, 84, 82]
Adaptability AI’s ability to adapt query strategy and learning based on human input. * Strategy Adaptation Effectiveness (improvement in query selection based on feedback), * Adaptability Score (composite metric of flexibility and responsiveness) [82, 85, 86]
Trust and Safety Trust in AI system, ensuring user safety. * Trust Score (level of user trust in AI system) [87], * Safety Incidents (number of safety-related issues) [88], * Error Reduction Rate (decrease in errors due to AI assistance) [2], * Confidence (level of user trust in AI recommendations) [89]
Task Allocation Complementarity AI handles data analysis and uncertainty estimation; Humans provide expertise and clarification. * Query Efficiency (amount of information gained per query), * Expertise Utilization (degree to which human expertise is leveraged) [78]
Flexibility Adjusting query strategy based on interaction dynamics. * Adaptability Score (same as in Interaction), * Dynamic Task Allocation (ability to adjust task distribution based on real-time needs) [23]
Efficiency Streamlining the learning process through targeted queries. * Model Improvement Rate (rate of improvement in AI performance) [59, 78], * Learning Curve (same as in Goals), * Resource Utilization (efficiency in using computational resources) [59]
Responsiveness AI’s ability to rapidly adjust to new inputs and feedback. * Response Time (time taken to respond to queries or feedback), * Real-time Performance (ability to handle time-sensitive tasks) [82, 90]
Collaborative Decision Making Joint decisions on data labeling and correction of AI’s model. * Impact of Corrections (measured improvement in system performance through human intervention) [91, 92], * Decision Effectiveness (impact of decisions on overall task success) [93, 94]
Continuous Learning AI’s ability to learn from human input and adjust its model. * Model Improvement Rate (same as in Efficiency), * Learning Curve (same as in Goals), * Knowledge Retention (ability to retain learned information over time) [95]
Mutual Support AI’s ability to provide support to humans in the learning process; Humans support by providing information. * Error Reduction Rate (decrease in errors due to AI assistance) [2], * Confidence (level of user trust in AI recommendations) [96], * Task Completion Time (with and without AI support) [97]
Robustness Ensuring AI system’s robustness and reliability. * Adversarial Robustness (resilience to adversarial attacks) [98], * System Reliability (overall system uptime and fault tolerance) [99]
Goals Factor
@@
The Goals factor ensures that the collaboration has a clear direction and that both human and AI efforts are aligned towards shared outcomes. This factor encompasses two key subfactors: Individual Goals, which pertain to the specific aims of each participant, and Collective Goals, which represent shared objectives that maximize the overall performance of the HAIC system.
Table 3 below summarizes the metrics used to evaluate these subfactors, providing concise definitions, measurement methods, and relevant references to support the evaluation of the Goals factor in Human-AI Collaboration (HAIC) systems.
Table 3: Summary of metrics for the Goals factor in Human-AI Collaboration.
Subfactor Metric Definition and Measurement
@@
Individual Goals Learning Curve Measures the rate of improvement over time, indicating the effectiveness of human teaching and AI learning. This is plotted by graphing AI performance (e.g., accuracy) over iterations or training samples. A steep learning curve suggests effective learning [76].
@@
Prediction Accuracy The proportion of correct predictions made by the AI model out of all predictions. This can be evaluated using metrics like accuracy, F1 score, or area under the ROC curve, depending on the task [80].
@@
Teaching Efficiency The gain in AI performance per unit of teaching effort by the human. This is measured by tracking improvements in AI accuracy or error rate after human input, as explored in studies such as Cakmak and Thomaz [79].
@@
Collective Goals Overall System Accuracy Calculated as a weighted average of performance metrics, reflecting the relative importance of each task in achieving the overall goal. This metric evaluates the general accuracy of the AI-human collaborative system, tailored to specific objectives [59].
@@
Objective Fulfillment Rate The percentage of predefined goals successfully achieved within the collaboration. Often used in project management contexts to track progress and adapted to assess HAIC success [81].
@@
The Interaction factor underscores the critical communication mechanisms through which humans and AI systems exchange information, provide mutual feedback, and adaptively respond to each other’s inputs. Key subfactors within this factor include Communication Methods, Feedback Mechanisms, Adaptability, and Trust and Safety. These subfactors capture the essential elements of interaction in Human-AI Collaboration (HAIC), ensuring seamless and effective communication, the incorporation of feedback, and adaptability to dynamic conditions.
Table 4 summarizes the metrics used to evaluate the Interaction factor, providing concise definitions and measurement methods for each subfactor.
Table 4: Summary of metrics for the Interaction factor in Human-AI Collaboration.
Subfactor Metric Definition and Measurement
@@
Communication Methods Clarity of Communication Assesses how understandable the communication is between humans and AI. This can be evaluated using qualitative methods like expert review or user feedback and quantitative measures like the proportion of correctly interpreted queries or time taken to resolve ambiguities [101, 86].
@@
Ease of Use Often measured through Likert scale questionnaires or user interviews to gauge the user’s subjective experience with the system.
Feedback Mechanisms Frequency of Feedback Monitors how often users provide feedback, helping developers assess engagement levels and identify areas for improvement [102].
@@
Feedback Quality Measures the relevance and specificity of feedback, either qualitatively through expert reviews or quantitatively by analyzing the correlation between feedback and subsequent AI performance improvements. Hein et al. [103] explore feedback acceptance and motivation when provided by AI versus human supervisors.
@@
Feedback Impact Quantifies the tangible effect of feedback on AI behavior or output quality, often measured by improvements in accuracy, precision, or other relevant metrics [104].
Adaptability Adaptability Score Combines measures of flexibility and responsiveness to assess how quickly and effectively the AI adapts to changes, as studied by El-Assady et al. [86].
@@
Strategy Adaptation Effectiveness Measures improvement in AI’s query selection based on human feedback, incorporating insights from semantic interaction research [85].
@@
Trust and Safety Trust Score Assesses user trust in the AI system through surveys and trust scales [105].
@@
Safety Incidents Tracks the number of safety-related issues encountered to evaluate the reliability and safety of the AI system.